# A Systematic Review of NLP for Dementia- Tasks, Datasets and Opportunities

### v1

@@ 3 Task Families
Figure 3 illustrates the contribution of various scientific communities to the research areas discussed above, including literature reviews and dataset introductions. While all communities participate in research on dementia detection as well as publish new datasets, the field of caregiver support has not yet been explored by the NLP and Speech communities. Additionally, we found no literature review similar to ours in NLP venues, highlighting a significant gap. We explore these and other gaps in the next section, pointing out opportunities for deeper, more rigorous research and innovative, exciting directions.
4 Gaps and Future Research
@@ 4.1 Data Variety
We begin by addressing the core of NLP research: the data. As noted in Section 3, the most popular datasets in the field are the Pitt corpus and its derivatives (ADReSS and ADReSSo), along with the CCC dataset (see additional details in Table 1). They are foundational and have significantly shaped the current state of NLP for dementia. However, like all data sources, they have limitations, including size constraints and demographic biases. For instance, the Pitt corpus contains more women than men and an imbalanced ratio of 35% healthy, 48% dementia, and 17% likely dementia participants. The ADReSS dataset was created as a balanced subset of Pitt, offering a smaller but evenly split sample by age, gender, and diagnosis Ševčík and Rusko (2022). Additionally, these datasets capture only the language of individuals with the access, ability, and willingness to participate in clinical trials. As a result, communities with less access to such trials, limited education, or intellectual and physical disabilities—factors known to impact cognitive assessments—are not represented Bruhn and Dammeyer (2018).
@@
60% of dementia patients are non-English speakers from low- and middle-income countries,888https://www.alzint.org/about/dementia-facts-figures/ making it crucial to gather data from these communities. Some studies try to bridge this data gap using translation methods and transfer learning for dementia detection in resource-limited languages (Drame et al., 2012; Fraser et al., 2019; Guo et al., 2020; Pérez-Toro et al., 2023, 2022; Kabir et al., 2023; Meng et al., 2023; Melistas et al., 2023). Nowenstein et al. (2024), who applied translation for dementia detection in Icelandic, caution that despite their promise, such methods need extensive validation due to varying grammatical complexities and dementia-related nuances across languages.
@@
Another concern regarding these classic datasets is their timeliness. For example, in the Pitt Corpus, published in 1994, the median participant age was 67, i.e., these individuals would be approximately 97 years old today. There is a significant difference between 67-year-olds now versus 30 years ago, including changes in technological proficiency and language use. This raises questions about whether models trained on these datasets can still provide accurate early diagnosis for today’s elderly population and whether the datasets could potentially be reproduced with updated protocols.
Another limitation of these datasets, and any dataset composed of transcribed audio, is their sensitivity to variations in automatic speech recognition (ASR) systems (Al Hanai et al., 2018; Balagopalan et al., 2019; Gómez-Zaragozá et al., 2023; Heitz et al., 2024). This issue is particularly critical for dementia-related data, where important cognitive indicators such as repetitions and silent pauses can be lost with certain ASR techniques.
Lastly, while these classic datasets significantly support research on tasks like detection and linguistic marker extraction, they offer no insight into caregivers, for example. They focus exclusively on patient data, overlooking other dementia stakeholders (Tahami Monfared et al., 2022), including the person conducting the cognitive assessment.
Contemporary
Let’s explore data sources beyond transcribed clinical trials. Publicly available data, like social media, has proven valuable for capturing the broad spectrum of communication across all dementia stakeholders (as discussed in Section 3). Additionally, some works creatively utilize non-dementia datasets like IMDB, leveraging its size and multilingual properties for translation methods and embedding extraction (Mirheidari et al., 2018; Li et al., 2019).
@@
An additional direction is freely available data of public figures diagnosed with dementia. Petti et al. (2023a, b) and Asllani and Mullen (2023) examine personal writings and spontaneous speech from well-known individuals with dementia, while Petti and Korhonen (2024) curated a corpus of a similar nature (see Appendix C). This data source is unique in spanning celebrities of different cultures and origins, different types of data sources (interviews, written texts, etc.), and different modalities (video, audio, etc.). Additionally, it allows for tracking these public figures throughout their disease journey, exploring their progression in an individualized way. Berisha et al. (2015) and Wang et al. (2020) emphasize this in their detailed case studies on the linguistic decline of President Ronald Reagan throughout his political career.
@@
Historically, ’NLP for dementia’ was synonymous with dementia detection, and annotated clinical records have proven well-suited for detection classifiers. However, to advance dementia research in all tasks and frontiers, NLP requires data that is abundant, diverse, and reflective of the disease’s complexity. Table 1 compares the traditional clinical datasets with the more contemporary ones we discussed, representing a small part of the 17 datasets we identified in our review (Table 2 in Appendix C). We hope these resources inspire researchers to pursue data from creative avenues, covering diverse dementia stakeholders, multiple languages, genders, cultures, and educational backgrounds, designing NLP solutions that are accessible to all.
@@ 4.2 Personalized Approaches
Every dementia patient experiences the disease’s progression uniquely, shaped by their individual characteristics and distinct personal language model. As each generation produces more accessible data, a 60-year-old today has likely generated significantly more text than a 60-year-old two decades ago. Given this wealth of textual data, can NLP facilitate personalized approaches for managing, tracking, or even diagnosing dementia by building individualized language models that monitor cognition on a personal level?
NLP-based digital twins for dementia patients represent a promising frontier in healthcare. These virtual models replicate a patient’s cognitive and behavioral patterns through their language, enabling personalized care by monitoring cognitive decline and adjusting treatment plans based on simulated behaviors Andargoli et al. (2024). A notable example is Hong et al. (2024), who developed digital twins by fine-tuning an LLM on longitudinal data from the I-CONECT clinical trial,999https://www.i-conect.org/ using these twins to evaluate a patient support chatbot.
An intriguing, yet unexplored, research direction involves creating personalized LLM companions modeled after loved ones, as cognitive stimuli from familiar individuals can positively impact dementia patients Woods et al. (2012). An LLM trained on family-curated data could engage patients with their personal history, family stories, and interests—such as favorite bands or recipes—potentially offering a non-clinical treatment. Given the vast amount of personal data generated today, this approach seems both feasible and effective. With these research directions, NLP has the potential to shape the future of personalized dementia treatments, both medical and alternative.
@@ 4.3 Trust and Scientific Rigor
A key contribution of our work lies in our ability to draw comparisons between the medical and technological communities, differing in various ways—one of which is scientific modeling. Medical researchers typically follow a hypothesis-driven process, designing studies and collecting data to test specific assumptions. In contrast, data scientists often retrospectively explore pre-collected data, searching for patterns without necessarily defining a specific hypothesis. This raises concerns for medical researchers: is the retrospective data relevant to future contexts (e.g., will a model trained on the Pitt corpus from 1994 capture the same signals in today’s dementia patients)? Can the model generalize well? Do the findings highlight correlation rather than causation? Indeed, the medical community, with its practical concerns, might be apprehensive about adopting solutions that, in their view, extract correlations from retrospectively collected, often small datasets, as this approach contrasts significantly with their scientific view of research. Their apprehension might be exacerbated by any lack of reporting on statistical significance or evaluation of robustness.
@@
To assess this, we annotated papers based on whether they report p-values or statistical significance tests, using it as a proxy for scientific rigor in NLP for dementia research.101010For literature reviews, statistical significance was marked as not relevant. Among all reviewed works, only 30% report statistical significance. While nearly half of NLP and medical papers include it, around 86% of works published in Speech venues —covering a significant portion of detection studies, particularly those using the popular ADReSS and ADReSSo datasets—do not.
Another crucial aspect of scientific modeling, alongside statistical significance, is robustness. One work addressing this is Novikova (2021), who examines the robustness of BERT-based dementia detection classifiers to data perturbations. Another example is Favaro et al. (2024), who investigates the consistency of linguistic features across multiple annotated corpora, identifying features that reliably differentiate between dementia and control groups while correlating with clinical scores.
A third aspect of scientific rigor is the adoption of explainability mechanisms in NLP-based solutions. This is particularly crucial for dementia detection algorithms that may eventually serve as real-world decision-support systems for doctors. Although some detection works prioritize explainable mechanisms (Ilias and Askounis, 2022a), such works remain relatively rare. The medical community is generally open to AI-based decision-support systems but demands transparency, understandably mistrusting "black-box" models (Adadi and Berrada, 2020). In fact, while the NLP community focuses on hard-to-decipher deep learning models—due to their higher detection accuracy—the medical community may be more inclined to adopt traditional machine learning classifiers given their transparency and interpretability.
@@
Ensuring scientific rigor is equally crucial—and challenging—for other task families. Works on caregiver support and patient assistance often lack a gold standard or clear objective, unlike the well-defined task of dementia detection. For example, evaluating LLMs that answer medical questions or simplify texts for dementia patients presents difficulties in terms of assessment and statistical rigor. Yet, understanding model outputs, assessing robustness, and clarifying motivations remain essential. Hence, developing benchmarks to evaluate these tools is a vital research direction.
Without transparent reporting of significance results, a focus on robustness and explainability, and clear evaluation benchmarks, NLP for dementia is at risk of remaining a tool limited to academic discussions, making its adoption into the medical community and other real-world applications significantly more challenging.
@@ 4.4 Application State-of-Mind
Alongside adopting rigorous methodologies and establishing trust, combating dementia in real-world settings requires NLP practitioners to embrace an application-focused mindset—’how can my work be applied in practice?’ A prominent example arises from our review of over 110 studies on NLP-driven dementia detection. Despite decades of promising results, we found no evidence in our research that any of these approaches have been integrated into real-world detection processes.111111We base this on publicly available data but acknowledge that efforts may exist beyond public knowledge
This raises the question: what is preventing the application of these promising NLP techniques in clinical settings? We hypothesize that trust issues (as discussed in the previous subsection) play a significant role, as it is simply too risky to adopt such tools into medical processes without rigorous testing across multiple datasets, contexts, and real-world scenarios. We believe that if researchers intentionally design their solutions with real-world applications in mind, outcomes may be different.
Applicative thinking goes beyond dementia detection—NLP tools for various tasks can be applied in practice with the right mindset, cross-community collaboration, and rigorous testing. For example, Topaz et al. (2020) developed a system to help home healthcare providers identify and cluster neuropsychiatric symptoms (e.g., anxiety, delusions), flagging dementia-related signs even in undiagnosed individuals. This system supports information transfer across healthcare settings and aids both professional and personal caregivers.
@@
Demand for dementia-related applications is evident from both business and technology perspectives. The dementia management and treatment market is expected to exceed a value of $36 billion USD by 2030,121212https://www.sphericalinsights.com/reports/dementia-drugs-market and independent technological solutions are quickly emerging. For instance, the ChatGPT ’app store’ now features over 30 dementia-related chatbots for cognitive stimulation or virtual companionship.131313Searched on https://chat.openai.com/, August 2024 People are taking initiative, leveraging LLMs to create these tools, and scientists can follow suit, using this framework to develop LLM-based playgrounds for large-scale applicative experiments. The need for practical NLP tools is undeniable, and we encourage the community to focus on solutions that are not only innovative but also truly applicable.
@@ 4.5 NLP Insights for Medicine
This work highlights the efforts of diverse scientific communities united by the common goal of advancing dementia research, and cross-disciplinary knowledge-sharing is imperative for success. For example, medical experts can collaborate with NLP practitioners when building datasets, while NLP researchers should work with medical professionals to develop robust and explainable solutions.
Knowledge transfer between the medical and technological communities should flow both ways. Currently, it is primarily the technological community (especially NLP) building on medical, linguistic, and cognitive knowledge in dementia research—for instance, using established linguistic markers and datasets originally collected through medical experiments like cognitive assessments. However, we argue that the other direction—where the medical community draws insights from NLP research—holds transformative potential.
@@
A specific example is the development of artificially degraded language models, which explore ’healthy’ vs. ’cognitively impaired’ models that replicate the linguistic behavior of dementia patients. Cohen and Pakhomov (2020) achieved this by fine-tuning an LSTM on the DementiaBank cohort, while Li et al. (2022) degraded different layers of GPT-2. Both studies impressively demonstrated dementia-like linguistic patterns in their models. Additionally, these works introduced perplexity as a useful metric for measuring cognitive decline, showing that degraded models exhibited significantly lower perplexity when exposed to dementia-related linguistic anomalies compared to ’healthy’ models. We believe that this research area can help us better understand the biological mechanisms of dementia and its pathologies. For example, an interesting direction would be to artificially degrade an LLM by mimicking protein deposits on neurons, potentially offering insights into the mechanisms of Alzheimer’s disease.
Loconte et al. (2023), on the other hand, question whether LLMs can truly enhance our understanding of the brain in the context of dementia. Testing ChatGPT with neuropsychological assessments for prefrontal functions, they find inconsistent results, showing these models still fall short of mimicking human cognition. This research sheds light on both cognitive mechanisms and how well neural networks model our brain’s complexity.
It’s not only the medical or biological community that could learn from NLP research. For example, the work of Li et al. (2024), which examines the parallels between the language of children and people with dementia, may advance research in cognitive sciences and linguistics. Using their artificially degraded language model mentioned above, they show that it is significantly less perplexed than GPT-2 when analyzing picture descriptions by young children and dementia patients. Looking at children above the age of 8, whose language aligns more closely with healthy adults, GPT-2 becomes less perplexed. This study highlights the potential for designing dementia-related biomarkers or novel cognitive assessments by learning from the language patterns of children.
We believe NLP has great potential for developing new cognitive assessment tests. Several studies suggest focusing on the dynamics between the interviewer and patient, rather than just the task at hand. They demonstrate that rare dialogue acts, such as clarification exchanges, and the interviewer’s role in the conversation can reveal cognitive signals that can potentially inform the creation of novel assessments (Farzana et al., 2020; Nasreen et al., 2021a; Farzana and Parde, 2022; Dawalatabad et al., 2022). Another example is Reeves et al. (2020), who introduced a novel NLP-empowered cognitive test based on narrative descriptions after watching video clips.
@@
Finally, NLP research may even prompt medical professionals to reconsider existing hypotheses about dementia risk factors. Zhou et al. (2019) analyzed a large cohort of text-rich health records and found a contradiction with contemporary studies that emphasize cardiovascular risks, such as high-fat, high-calorie, and high-carbohydrate diets, as causes of dementia. Surprisingly, these factors were rarely mentioned in patients’ clinical records. Instead, they found significant associations with deficiencies in potassium, calcium, and other nutrients, mentioned in over 25% of dementia patients’ records. However, actual laboratory results for nutrient levels were rarely documented. This study highlights NLP’s potential to challenge existing hypotheses, and expose gaps in medical procedures.

### v2

@@ 3 Task Families
Figure 3 shows the contributions of different scientific communities to the discussed task families, including literature reviews and dataset introductions. While all communities engage in dementia detection and dataset publication, caregiver support remains unexplored in NLP and Speech research. Additionally, we found no literature review similar to ours in NLP venues, highlighting a significant gap. The next section explores these and other gaps, highlighting opportunities for deeper, more rigorous research, as well as ethical dilemmas and innovative directions.
4 Gaps and Opportunities
@@ 4.1 Data Variety
We begin our discussion of gaps by addressing a core aspect of NLP research: the data. As noted in Section 3, the most popular datasets in the field are the Pitt corpus and its derivatives (ADReSS and ADReSSo), along with the CCC dataset (see additional details in Table 1). These datasets are foundational and have significantly shaped the current state of NLP for dementia. However, like all data sources, they have limitations, including size constraints and demographic biases. For instance, the Pitt corpus contains more women than men, as well as an imbalanced ratio of 35% healthy, 48% dementia, and 17% likely dementia participants. The ADReSS dataset was created as a balanced subset of Pitt, offering a smaller but evenly split sample by age, gender, and diagnosis Ševčík and Rusko (2022). Additionally, these datasets capture only the language of individuals with the access, ability, and willingness to participate in clinical trials. As a result, communities with less access to such trials, limited education, or intellectual and physical disabilities - factors known to impact cognitive assessments - are not represented Bruhn and Dammeyer (2018).
@@
60% of dementia patients live in low- and middle-income countries,666https://www.alzint.org/about/dementia-facts-figures/ making it crucial to gather data multilingual data from such communities. Some studies try to bridge this gap using translation, transfer learning, or cross-lingual statistical methods for resource-limited languages (Drame et al., 2012; Fraser et al., 2019; Lindsay et al., 2021; Guo et al., 2020; Pérez-Toro et al., 2023, 2022; Kabir et al., 2023; Meng et al., 2023; Melistas et al., 2023). One such example is Nowenstein et al. (2024), who applied translation for dementia detection in Icelandic. They caution that despite their promise, such methods need extensive validation due to varying grammatical complexities and dementia-related nuances across languages.
@@
Another concern with these classic datasets is their timeliness. For instance, the Pitt Corpus, published in 1994, had a median participant age of 67. There is a significant difference between 67-year-olds now versus 30 years ago, including changes in technological proficiency and language use. This raises questions about whether models trained on these datasets are effective for detection in today’s elderly population, and whether these datasets may be reproduced with updated protocols.
Additionally, these classic datasets were manually transcribed. As such, they contain clean, standardized transcriptions with expert annotations of linguistic nuances such as disfluencies, pauses, and correction. However, manual transcription and annotation are neither scalable nor cost-effective. In real-world applications, texts will therefore likely be transcribed using automatic speech recognition (ASR), which produces noisier outputs and often fails to capture critical linguistic markers of cognitive decline (Al Hanai et al., 2018; Balagopalan et al., 2019; Gómez-Zaragozá et al., 2023; Heitz et al., 2024). As a result, models trained on these clean, expert-transcribed datasets may struggle to generalize to real-world settings.
A final limitation of these classic datasets stems from the dementia stakeholder they represent- namely, the patient. Other humans involved, such as caregivers, family members, and even the individuals conducting the cognitive assessments Tahami Monfared et al. (2022), are not represented. Given the data-driven nature of NLP, this may naturally cause practitioners to focus on patients and inadvertently constrain the scope of research on all other dementia stakeholders.
@@
Contemporary
We now turn to data sources beyond clinical trials. Social media posts have proven valuable for capturing the broad spectrum of communication across various dementia stakeholders (discussed in Section 3). Some studies creatively used non dementia-specific datasets, like IMDB, leveraging their size and multilinguality or translation or embedding extraction for dementia-related tasks (Mirheidari et al., 2018; Li et al., 2019).
An additional direction is freely available data of public figures diagnosed with dementia. Petti et al. (2023a, b) and Asllani and Mullen (2023) examine personal writings and spontaneous speech from well-known individuals with dementia, while Petti and Korhonen (2024) curated a corpus of a similar nature (mentioned in Table C). This data source is unique in spanning celebrities of different cultures and origins, various types of data sources (interviews, written texts, etc.), and multiple modalities (video, audio, etc.). Additionally, it allows for tracking these public figures throughout their disease journey, exploring their individual progression. Berisha et al. (2015) and Wang et al. (2020a) emphasize this in their detailed case studies on the linguistic decline of President Ronald Reagan throughout his political career.
@@
Historically, ‘NLP for dementia’ was synonymous with dementia detection, and the classic datasets containing annotated clinical records have proven well-suited for detection classifiers. However, to advance dementia research in all tasks and frontiers, NLP requires data that is abundant, diverse, and reflective of the disease’s complexity. Table 1 compares the traditional clinical datasets with the more contemporary ones we discussed, representing a small part of the 17 datasets we identified in our review (Table 2 in Appendix C). We hope these resources inspire researchers to pursue data from creative sources, covering diverse dementia stakeholders, multiple languages, genders, cultures, and educational backgrounds, designing NLP solutions that are accessible to all.
@@ 4.2 Personalized Approaches
Every dementia patient experiences the disease progression uniquely, shaped by their individual characteristics and ‘distinct personal language model’. As each generation produces more accessible data, a 60-year-old today has likely generated significantly more text than a 60-year-old two decades ago. Given this wealth of textual data sourced from older adults, NLP may facilitate personalized approaches for managing, tracking, or even diagnosing dementia by building individualized language models that monitor cognition on a personal level.
NLP-based digital twins for dementia patients represent a promising frontier in healthcare. These virtual models replicate a patient’s cognitive and behavioral patterns through their language, enabling personalized care by monitoring cognitive decline and adjusting treatment plans based on simulated behaviors Andargoli et al. (2024). A notable example is Hong et al. (2024), who developed digital twins by fine-tuning an LLM on longitudinal data from the I-CONECT clinical trial,777https://www.i-conect.org/ using these twins to evaluate a patient support chatbot.
An intriguing, yet unexplored, research direction involves creating personalized LLM companions modeled after loved ones, as cognitive stimuli from familiar individuals can positively impact dementia patients Woods et al. (2012). An LLM trained on family-curated data could engage patients with their personal history, family stories, and interests - such as favorite bands or recipes - potentially offering a non-clinical treatment. Given the vast amount of personal data generated today, this approach seems both feasible and effective. With these research directions, NLP has the potential to shape the future of personalized dementia treatments, both medical and alternative.
@@ 4.3 Trust and Scientific Rigor
A key contribution of our work is drawing comparisons between the medical and technological communities, particularly in scientific modeling. Medical research typically begins with a hypothesis, followed by study design and data collection for analysis. In contrast, data scientists often start with pre-collected data, retrospectively identifying patterns without necessarily defining a specific hypothesis. This conceptual gap may be critical in medical contexts. Is retrospective data relevant to future cases (e.g., will a model trained on the 1994 Pitt corpus capture the same signals in today’s dementia patients)? Do the results indicate correlation rather than causation? The medical community may be hesitant to adopt solutions they perceive as correlations derived from retrospectively collected, often small datasets El-Sappagh et al. (2023).
@@
Hesitance to adopt NLP-based solutions may grow when studies do not report statistical significance or assess robustness. We manually annotated each paper in our cohort, excluding literature reviews, to indicate whether statistical tests or p-values were reported. Among all reviewed papers, only 33% report statistical significance. While about half of NLP and medical papers include it, nearly 86% of Speech publications - covering a significant portion of detection studies, particularly those using the popular ADReSS and ADReSSo datasets - do not (see Figure 4).
Another crucial aspect of scientific modeling is robustness. We found few studies in our cohort which explicitly addressed this topic or evaluated the robustness of their models. Two notable examples are Novikova (2021), who examined BERT-based dementia detection under data perturbations, and Favaro et al. (2024), who assessed the consistency of linguistic features across corpora.
Explainability, a key pillar of scientific rigor, remains rare in dementia detection, with few exceptions like Ilias and Askounis (2022a). A disconnect exists between the NLP community, which favors deep learning for higher accuracy, and the medical community, which prioritizes transparency and tends to distrust "black-box" models (Adadi and Berrada, 2020). As a result, clinicians often prefer traditional, inherently interpretable models. Another gap exists as NLP practitioners’ favor feature-attribution methods, explaining prediction by assigning importance scores to individual tokens or features (Karlekar et al., 2018; Ilias and Askounis, 2022a; Vimbi et al., 2024). While straightforward for researchers, these methods often produce overwhelmingly low-level explanations (e.g., assigning an importance score to every token in a lengthy transcript), potentially confusing both clinicians and patients, especially those with cognitive impairments. Addressing these challenges requires cross-disciplinary collaboration to develop interpretability methods that are both scientifically rigorous and clinically meaningful. We encourage researchers to consult surveys such as Stiglic et al. (2020); Calderon and Reichart (2024); Viswan et al. (2024), which outline NLP interpretability paradigms applicable across various algorithms.
@@
Ensuring scientific rigor is just as crucial - and equally challenging - for tasks beyond dementia detection. Studies on caregiver support and patient assistance often lack a gold standard or clear objective. For example, evaluating LLMs that answer medical questions or simplify texts for dementia patients poses challenges in assessment and statistical rigor. Developing benchmarks to evaluate these tools is therefore a vital research direction.
@@ 4.4 Application State-of-Mind
To combat dementia in real-world settings, NLP practitioners must not only adopt rigorous methods and build trust- they should also consider the applicability of their work, and ask themselves - ‘how can my work be applied in practice?’ Dementia detection exemplifies this gap- despite decades of promising results and hundreds of surveyed papers, our research found no evidence that these approaches have been integrated into real-world detection processes.888Based on publicly available data. What hinders the clinical adoption of such promising NLP techniques? First, we believe that trust issues (as discussed earlier) are a major factor, as integrating these tools into medical workflows is too risky without rigorous testing across diverse datasets and real-world scenarios. However, they are not the only barrier.
One applicative challenge in dementia detection lies in the reliance on spoken responses, which must be transcribed in real time or post-recording. Clinicians often cannot record assessments due to privacy concerns and technological constraints, and even if recordings were conducted regularly, transcription would remain a bottleneck. Manual transcription is costly and time-consuming, making automatic speech recognition (ASR) the most practical alternative. Many studies have focused on improving ASR for dementia-related tasks - aligning transcriptions (Lehr et al., 2012, 2013), tailoring ASR to specific dementia datasets (Abulimiti et al., 2020), and fine-tuning models like Whisper (Li and Zhang, 2024; Radford et al., 2023).Yet, ASR systems still introduce errors, miss key markers like disfluencies and repetition, and obscure the distinction between patient incoherence from transcription errors Li et al. (2024b). NLP practitioners must consider the entire pipeline, not just their models. If solutions depend on error-prone ASR, they risk adding complexity rather than reducing burden.
Applicative thinking extends beyond dementia detection, relevant to all dementia-related tasks. Topaz et al. (2020), Zolnoori et al. (2023) and Ryvicker et al. (2024), who tailor their NLP applications to the home healthcare setting, help professional caretakers in identifying, clustering, and analyzing dementia symptoms and behaviors (e.g., anxiety, delusions, hoarding). These systems were designed for real-world deployment, aiming to flag dementia-related signals in undiagnosed individuals and facilitate information transfer across healthcare settings, supporting both professional and personal caregivers. Casu et al. (2024) present another form of applicative consideration, designing LLMs for dementia-related tasks and optimizing them for local fine-tuning to ensure data safety - an approach that might be more acceptable to the healthcare community compared to cloud-based training.
@@
Demand for dementia-related applications is evident from a business perspectives as well. The dementia treatment market is projected to surpass $36 billion by 2030, with independent technological solutions rapidly emerging.999https://www.sphericalinsights.com/reports/dementia-drugs-market For instance, the ChatGPT ‘app store’ now features over 30 dementia-related chatbots for cognitive stimulation or virtual companionship.101010Searched on https://chat.openai.com/, August 2024 People are taking initiative, leveraging LLMs to create these tools, and scientists can follow suit, using this framework to develop LLM-based playgrounds for large-scale applicative experiments. Such applications may also empower older adults to administer self-assesment and monitoring, as shown by Skirrow et al. (2022).
To summarize, the need for practical NLP solutions is undeniable, and we encourage the community to focus on solutions that are not only innovative but also truly applicable in real-life settings.
@@ 4.5 NLP Insights for Medicine
With many scientific communities advancing dementia research, cross-disciplinary collaboration is key. While NLP and other technological fields draw from medical insights and data, we argue that the reverse - integrating NLP insights into medical science and practice - holds immense potential.
@@
One example is the concept of artificially degraded language models, which explore ‘healthy’ vs. ‘cognitively impaired’ models replicating the linguistic behavior of dementia patients. Cohen and Pakhomov (2020) achieved this by fine-tuning an LSTM on the DementiaBank cohort, while Li et al. (2022a) degraded different layers of GPT-2. Both approaches yielded models that impressively demonstrated dementia-like linguistic patterns. The two studies utilized perplexity (a derivative of cross-entropy) as a measure of cognitive decline, a concept introduced in Roark et al. (2011). The degraded models exhibit significantly lower perplexity when exposed to dementia-related linguistic anomalies compared to ‘healthy’ models. We believe that this research area might shed light on the biological mechanisms of dementia and its pathologies- for example, degrading an LLM by mimicking protein deposits on its neurons might offer insights into the mechanisms of Alzheimer’s.
Li et al. (2024a) use a degraded model to explore similarities between children’s language and that of dementia patients. Their model shows lower perplexity than GPT-2 on picture descriptions by young children and dementia patients, and significantly higher perplexity on descriptions by children over 8, whose language aligns with that of with healthy adults. This work suggests NLP can aid in developing dementia biomarkers or cognitive assessments inspired by children’s language.
NLP can also help craft novel cognitive assessments that examine more than the patient’s performance of the task at hand. Farzana et al. (2020); Nasreen et al. (2021a); Farzana and Parde (2022) and Dawalatabad et al. (2022) analyze interaction dynamics between patients and interviewers, demonstrating that rare dialogue acts, such as clarification exchanges and interviewer utterances, can signal cognitive variations. Reeves et al. (2020) propose an alternative assessment, focusing on narrative descriptions following video observations.
Another research direction can illuminate patients’ lifestyle factors and experiences for the medical community. Studies such as Yi et al. (2020); Shen et al. (2022); Wu et al. (2023) and Zhou et al. (2018) focus on extracting patients’ lifestyle factors from clinical records and social worker notes. They identify, for example, whether patients experience sleep deprivation, suffer from social isolation, or report substance abuse. Such NLP-based automated approaches can provide insights into a patient’s unique experience of the disease and even allow for cognitive profiling Upadhyay et al. (2022), potentially allowing clinicians to tailor interventions that improve individualized patient outcomes.
@@
Finally, NLP can prompt clinicians to reconsider existing hypotheses about dementia risk factors. Zhou et al. (2019, 2021) conducted extensive analyses of health records, uncovering contradictions with prevailing medical assumptions. While many medical studies emphasize cardiovascular risks - such as high-fat, high-calorie, and high-carbohydrate diets - these factors were rarely noted in dementia patient records. Instead, deficiencies in potassium, calcium, and other nutrients, as well as frequent mentions of dietary supplements (or their absence), appeared in over 25% of dementia patients’ records. Actual lab results for nutrient levels, however, were rarely documented, a result which challenges existing hypotheses, and exposes potential gaps in medical procedures.
5 Directions for Future Work
Previous sections outlined the state of NLP for dementia and identified gaps and opportunities. Along the way, we highlighted open challenges that struck us as meaningful, practical, and yet to be fully explored (see Figure 5). We now distill these into concrete takeaways for researchers.
In Section 4.2 we discussed personalized LLMs as a prominent direction. Namely, digital twins - personalized language models mimicking a patient’s linguistic patterns - could help simulate disease progression, predict responses to clinical trials, or generate patient-specific data. Alternatively, LLMs based on loved ones - i.e., fine-tuned on texts from a patient’s family and friends - could enable truly personalized emotional and cognitive care. However, to ensure rigorous evaluation and adoption, benchmarks for personalized LLMs are a necessity. Researchers and domain experts must collaborate to create datasets, define ‘desired’ and ‘undesired’ behaviors, and establish standards for reliability, factuality, emotional capabilities, biases, risks, and ethical considerations (Section 6).
From a data perspective, we urge researchers to explore synthetic data augmentation- for example, making transcribed clinical assessments more representative of populations unable to attend (as discussed in Section 4.1). Enhancing data with culturally diverse linguistic patterns, such as accents and regional slang, can help include these individuals without requiring direct trial participation.
Researchers should also prioritize multilingual, multimodal datasets to ensure dementia research worldwide reflects true linguistic, demographic, and cultural diversity. All stakeholders - patients, caregivers, clinicians, maybe even dementia-related content creators - should be represented. Another important direction is reproducing classic protocols like Pitt to interview older adults and collect data. This could evaluate their relevance, track generational language shifts, and test whether models trained on data from such protocols can generalize to today’s population of older adults.
From a medical perspective, we believe that artificially degraded language models (discussed in Section 4.5) are the next frontier. Researchers should explore this by modeling pathologies with LLMs - disrupting specific model layers to mimic frontotemporal dementia or mimicking protein deposits on neurons to simulate Alzheimer’s. They can then analyze whether linguistic nuances emerge (e.g., does simulated frontotemporal dementia lead to more aggressive language?) and explore what might improve the LLM’s ‘condition.’
Finally, for NLP to truly support early detection, dementia must be diagnosed as early as possible. While it’s tempting to showcase high accuracy on Alzheimer’s cases, this has limited clinical value since (a) clinicians can usually diagnose Alzheimer’s easily, and (b) few interventions exist at that stage. Researchers must therefore focus on MCI- develop algorithms attuned to the subtlest linguistic shifts and prioritize datasets with sufficient MCI cases - or even create MCI-specific datasets, such as texts from the recently diagnosed.
These research paths demand careful considerations, especially given the ethical complexities explored in the next section.
6 Ethical Discussion
Ethical dilemmas regarding AI and healthcare have been widely examined (Rigby, 2019; Gerke et al., 2020; Keskinbora, 2019; Harrer, 2023). However, in this section, we highlight those specific to NLP and cognitive decline. One example is collecting data from the cognitively impaired- an ethically delicate task, as these individuals often cannot provide informed consent, raising questions about autonomy and data use. Fairness and diversity is another challenge, as NLP systems may misinterpret silent pauses, repetitions, or corrections, mistaking them for cognitive decline when they actually stem from non-native speech or even simply from stress.
The classic precision-recall trade-off also takes on new ethical weight in dementia research, particularly in detection. For treatable conditions like certain cancers, prioritizing recall to err on the side of caution is often justified. But for untreatable, progressive illnesses like dementia, false positives can be deeply harmful. A misdiagnosis, especially at a younger age, can lead to severe emotional distress, depression, and even suicidal thoughts (Lyketsos and Lee, 2003; Naismith et al., 2022). Misdiagnosed individuals may even be prescribed newly emerging drugs and treatments, exposing them to serious side effects like brain edema and hemorrhages- despite not needing the medication.
Another challenge may arise from using LLMs to generate patient speech. LLM hallucinations - non-factual outputs Smith et al. (2023) - and confabulations - false memories without intent to deceive Brown et al. (2017) - are problematic for many NLP tasks but are also key characteristics of dementia patients. How does one distinguish between LLM fabrications and authentic simulations of cognitive impairment? This raises both technical and ethical concerns: while some generated patterns (e.g., excessive empty speech or extreme hallucinations) may enhance accuracy, they can also introduce bias and reduce interpretability. Misleading data risks distorting our understanding of dementia, its cognition, and lived experience. Establishing benchmarks and ethical guidelines for synthetic data in this domain is essential.
Finally, researchers must ensure that models communicate predictions responsibly and safely. Dementia patients, often experiencing impaired judgment, may overshare sensitive information, act on unsafe advice, or follow suggestions with unintended consequences. For instance, if a chatbot recommends taking a walk to feel better, a patient might do so and become lost- an entirely plausible scenario. Addressing these ethical challenges demands cross-disciplinary collaboration, rigorous risk assessment, robust evaluation, and a strong sense of responsibility and accountability.
7 Limitations
Like all surveys, our work represents a snapshot in time. With NLP advancing extremely rapidly, new relevant studies are highly likely emerge after this review’s publication. Additional limitations stem from deliberate choices to maintain a manageable scope and focused analysis. For instance, we relied on relatively broad dementia-related terms rather than expanding to all specific pathologies (e.g., Lewy Body, Vascular) or other potentially related medical conditions (e.g., Parkinson’s disease). Additionally, we focused on studies using data that is fully or partially in English, thus excluding research conducted solely in other languages. Finally, we confined our survey to peer-reviewed articles, acknowledging that in NLP, innovation often originates in open-source publications. Despite these limitations, we believe our review provides comprehensive coverage of the NLP for Dementia domain and provides a solid starting point for researchers looking to advance this field.